### v2

# On the Self-awareness of Large Reasoning Models’ Capability Boundaries

Abstract
Large Reasoning Models (LRMs) have shown impressive performance on complex reasoning tasks such as mathematics, yet they also display misbehaviors that expose their limitations. In particular, when faced with hard questions, LRMs often engage in unproductive reasoning until context limit, producing wrong answers while wasting substantial computation. This phenomenon reflects a fundamental issue: current answering paradigms overlook the relationship between questions and LRMs’ capability boundaries. In this paper, we investigate whether LRMs possess self-awareness of capability boundaries. We begin by an observation that LRMs may know what they cannot solve through expressed reasoning confidence. For black-box models, we find that reasoning expressions reveal boundary signals, with accelerated growing confidence trajectory for solvable problems but convergent uncertainty trajectory for unsolvable ones. For white-box models, we show that hidden states of the last input token encode boundary information, with solvable and unsolvable problems linearly separable even before reasoning begins. Building on these findings, we propose two simple yet effective optimization strategies: reasoning expression monitoring and hidden states monitoring. Experiments demonstrate that these boundary-aware strategies enable LRMs to avoid unproductive reasoning without sacrificing accuracy, significantly improving reliability and efficiency by cutting token usage up to – . We open-source at https://github.com/qingjiesjtu/CapBound.
1 Introduction
Large Reasoning Models (LRMs) have demonstrated remarkable capabilities on complex reasoning tasks such as mathematics (Guo et al., 2025; Jaech et al., 2024; Ahn et al., 2024). However, they also exhibit a range of misbehaviors that reveal the limitations of their capabilities (Kalai et al., 2025; Yao et al., 2025b; Sun et al., 2025; Zhang et al., 2024). In particular, when confronted with hard questions, LRMs may fall into repetitive looping until the context length is exhausted (Zhou et al., 2025; Chen et al., 2025b). Such a response paradigm not only fails to produce a correct answer but also wastes computational resources (Peng et al., 2025; Pu et al., 2025; Chen et al., 2024b), leaving users waiting a long time with no meaningful outcome.
This phenomenon highlights a fundamental issue: current answering paradigms of LRMs overlook the relationship between questions and their capability boundaries. For questions beyond their capability boundary, LRMs still struggle to produce a wrong answer, often engaging in unproductive long reasoning manifested as repetitive looping or error accumulation (Yao et al., 2025a; Mukherjee et al., 2025) (shown in Appendix E), which undermines both response efficiency and reliability. Therefore, it is crucial to identify questions beyond capability boundaries, and to explicitly acknowledge the inability to solve them. This importance has also been recognized by the broader AI community, as exemplified by the praise for GPT-5’s ability to appropriately say “I don’t know.”111https://x.com/koltregaskes/status/1957474061153436094
Existing approaches to exploring capability boundaries are mainly conducted from an external perspective, for example, by constructing increasingly challenging benchmarks such as Humanity’s Last Exam (HLE) (Phan et al., 2025), or by designing real-time routers that quickly estimate question difficulty and assign harder questions to more capable models (OpenAI, 2025). However, little work has investigated capability boundaries from the model’s internal perspective, namely self-awareness of capability boundaries, which is particularly meaningful for understanding reasoning mechanism and for further improving reliability and efficiency.
Thus, this paper focuses on answering the question: “Do LRMs know their capability boundaries?” Figure 1 shows the overview of our work. We begin with the observation that LRMs often exhibit awareness of what they cannot solve (Section 3), and we provide evidence for such self-awareness for both black-box and white-box scenarios. For black-box models, we show that reasoning expressions contain clear signals of capability boundaries (Section 4). For white-box models, we find that hidden states encode information that also delineates these boundaries (Section 5). Taken together, these findings suggest that LRMs can assess, on the fly, whether a given question is beyond capability boundary, thereby enabling more efficient and reliable reasoning (Section 6). We summarize our contributions as follows.
We observe indications of LRMs’ self-awareness of capability boundaries by comparing their reasoning confidence, articulated in anthropomorphic tones, with the correctness of their final answers. Notably, LRMs know what they cannot solve.
From a black-box perspective, we reveal LRMs’ capability boundaries through reasoning trajectories. Remarkably, their reasoning shows an accelerated growing confidence trajectory for questions within the boundary, but a convergent uncertainty trajectory for questions beyond it.
From a white-box perspective, we reveal LRMs’ capability boundaries through the hidden states of the last input token. Remarkably, solvable and unsolvable questions are almost linearly separable in the hidden space, even before the reasoning begins.
To leverage capability boundaries during reasoning, we design two simple yet effective optimization strategies: Reasoning Expression Monitoring (black-box) and Hidden States Monitoring (white-box). Experiments show that these boundary-aware strategies substantially improve efficiency and reliability without sacrificing accuracy, reducing token usage up to – .
2 Related Work
From knowledge boundary to capability boundary. Recent studies have examined the knowledge boundaries of large language models (LLMs), focusing on the limits of factual knowledge encoded in parameters (Yin et al., 2024; Li et al., 2024). For instance, Wen et al. (2024) investigate LLMs’ perception of knowledge boundary through semi-open-ended question answering. Deng et al. (2025) unveil knowledge boundary of LLMs for trustworthy information access. Kale & Nadadur (2025) evaluate LLMs’ self-knowledge via consistency. These works primarily emphasize factual coverage, with retrieval-augmented generation (RAG) being a common solution to extend the accessible knowledge space (Gao et al., 2023; Lewis et al., 2020).
By contrast, the capability boundary concerns what reasoning tasks LLMs can reliably solve (Chen et al., 2025a; Xue et al., 2025; Chen et al., 2024a). This notion is broader yet rarely formalized. Existing efforts typically approximate it externally: either by constructing more challenging benchmarks such as Humanity’s Last Exam (HLE) (Phan et al., 2025), or by designing routing mechanisms such as the GPT-5 router that delegate harder questions to more capable models (OpenAI, 2025). In contrast, this paper investigates capability boundaries from the internal perspective of LLMs, focusing on their potential self-awareness.
Interpretability of reasoning expressions and hidden states. Interpretability aims to uncover how LLMs understand and execute tasks, providing insights that enhance their reliability and efficiency. Recent studies have approached from both reasoning expressions and hidden states. On the reasoning expressions side, Guo et al. (2025) highlight the anthropomorphic cues, “aha moments”, as a key role in reasoning. Qian et al. (2025) tracks mutual information along the generation and finds sharp surges at specific thinking tokens (e.g., “Hmm”, “Wait”, “Therefore”). Bogdan et al. (2025) frame sentence-level importance as a counterfactual influence problem, showing that only a few critical utterances steer the reasoning trajectory.
On the hidden states side, Wendler et al. (2024) probe multilingual Llama models and reveal English as the pivot language in the internal structure. Fan et al. (2024) demonstrate that certain intermediate layers already encode sufficient predictive features, implying redundancy in the full forward pass. Zhang et al. (2025) train linear probes on hidden states to verify intermediate answers, anticipating future correctness. Our work builds on both perspectives by analyzing reasoning expressions and hidden states to reveal self-awareness of capability boundaries.
Efficient reasoning: from solvable questions to unsolvable questions. A range of recent works examine the phenomenon of overthinking (Sui et al., 2025). Chen et al. (2024b) show that LRMs often allocate excessive computation to easy questions, while Zhang et al. (2024) report that multi-turn self-correction can lead to unnecessary cost. To mitigate such inefficiency, Wang et al. (2025) demonstrate that suppressing frequent reflective tokens such as “Wait”, “Hmm”, or “Alternatively” improves reasoning efficiency without harming accuracy. Similarly, Yang et al. (2025a) introduce mechanisms to decide during reasoning whether further steps are unnecessary, allowing the model to terminate early and output the final answer directly.
However, these works target the case where LRMs can solve a question but waste resources by reasoning too long. In contrast, our work addresses the complementary scenario: when a question is beyond the model’s capability boundary, LRMs still continue reasoning unproductively despite having no chance to get a correct answer.
3 Large Reasoning Models Know What They Cannot Solve
We begin our exploration by evaluating commonly used LRMs including GPT-oss-20B (Agarwal et al., 2025), DeepSeek-R1-0528-Qwen3-8B (abbreviated as R1-Distill-Qwen3-8B) (Guo et al., 2025), DeepSeek-R1-Distill-Qwen-32B (abbreviated as R1-Distill-Qwen-32B) (Guo et al., 2025), and QwQ-32B (Qwen Team, 2024) on several standard mathematical benchmarks, including AIME’25 (MAA, 2025), AIME’24 (MAA, 2024), HMMT (HMMT, 2025), and AMC’23 (MAA, 2023) (see Appendix A for detailed descriptions of datasets and models).
Inspired by works which monitors LRMs’ reasoning (Baker et al., 2025; He et al., 2025; Luo et al., 2024), we examine the reasoning processes (e.g., the segments wrapped within <think> in DeepSeek-like models) to look for signs of self-awareness regarding whether the input question can be correctly solved. As illustrated in Figure 2, we observe that such signals, expressed as anthropomorphic tones (Guo et al., 2025; Yang et al., 2025b), emerge throughout the reasoning process, and become especially evident in its final stages. Toward the end of reasoning, models often produce explicit statements that reflect their confidence level. Strikingly, these expressed beliefs tend to align with the correctness of their final answers. For example, when the reasoning contains more confident expressions such as “So, no mistake” or “I think it’s correct”, the final answers are typically correct. Conversely, when uncertain expressions like “My initial assumption is wrong” or “I’m not 100% sure” appear, the final answers are incorrect.
This observation motivates us to quantify the alignment between expressed belief and answer correctness, which we formalize through a pair of metrics:
| Can % | (1) | |||
| Cannot % |
Specifically, Can % is the proportion of questions for which the model expresses more confidence during reasoning and indeed produces a correct answer, while Cannot % is the proportion of questions for which the model expresses more uncertainty and indeed fails to provide the correct answer.
| Can % | Cannot % | |
| GPT-oss-20B | ||
| R1-Distill-Qwen3-8B | ||
| R1-Distill-Qwen-32B | ||
| QwQ-32B |
Table 1 shows the results. We observe that LRMs’ expressed beliefs align with the correctness of their final answers. In particular, for questions where LRMs express more uncertainty during reasoning, the final answers are consistently incorrect (Cannot % = 100). This phenomenon is observed across multiple models, suggesting that LRMs are aware of what they cannot solve. Taken together, the reasoning expressions indicate a form of self-awareness of capability boundary.
4 Capability Boundary Revealed by Reasoning Expressions
Motivated by the observations in the previous section, we hypothesize that it is feasible to reveal LRMs’ capability boundaries through their reasoning expressions. Therefore, we systematically investigate the dynamics of confident and uncertain expressions throughout the reasoning process.
We first construct two sets of expressions, confident and uncertain expressions, that frequently appear in the reasoning process (the full list is provided in Appendix C). These anthropomorphic expressions are commonly observed and have been recognized in prior work as critical for reasoning (Yang et al., 2025b; Guo et al., 2025; Qian et al., 2025). Second, we divide the reasoning process into multiple stages (e.g., 50 stages) and compute the density of confident and uncertain expressions at each stage. Finally, concatenating these values yields the expression density trajectories, denoted as for confident expressions and for uncertain expressions. These trajectories capture how the density of each type of expression evolves over the course of reasoning.
We plot and in Figure 3, separately for cases with correct and wrong final answers. We observe that LRMs exhibit clear differences in their reasoning expression trajectories depending on final answer correctness. When the final answer is correct, the confident trajectory lies above the uncertain trajectory and shows an accelerating increasing trend. In contrast, when the final answer is incorrect, the uncertain trajectory dominates over the confident trajectory and tends to converge. Based on this distinction, the capability boundary can be defined as a linearly increasing trajectory (blue line in Figure 3). A concave trajectory indicates that the question lies within the LRM’s capability boundary, while a convex trajectory indicates that the question lies beyond it.
To formally characterize these patterns, we design two indicators as criteria for determining whether a question is within or beyond the capability boundary:
Confidence Differential (ConfDiff) measures the sign of gap between uncertain and confident trajectory along the reasoning process. Specifically, we accumulate the sign up to a given reasoning stage percentage , and treat the question as beyond the capability boundary if the cumulative value exceeds a threshold :
(2) -
Confidence Curvature (ConfCurv) computes the second derivative of the expression density trajectory to capture its convexity (Boyd & Vandenberghe, 2004). A concave trajectory corresponds to questions within the capability boundary, while a convex trajectory indicates questions beyond it. Also, we accumulate the sign of convexity up to a reasoning stage percentage and treat the question as beyond the capability boundary if the value exceeds a threshold :
Figure 4 shows that the two proposed indicators can reliably distinguish unsolvable from solvable questions, achieving high accuracy across different LRMs throughout the reasoning stage. Between the two indicators, ConfDiff is more reliable as it achieves higher accuracy and exhibits smaller fluctuations during the reasoning process. Moreover, the high accuracy emerges early, e.g., as early as 2% for GPT-oss-20B. Interestingly, the gradually declining yellow curve of this model suggests that boundary signals are more salient at the beginning than at the end of reasoning. This is likely due to repetitive looping and error accumulation in later stages (Yao et al., 2025a; Mukherjee et al., 2025). Therefore, we demonstrate that the capability boundary can be revealed through reasoning expressions, even at an early stage.
5 Capability Boundary Revealed by Hidden States
Following the observation that capability boundaries can be revealed at an early stage of reasoning, it is worth examining whether such signals may emerge even earlier, potentially before the model begins reasoning. Therefore, we investigate whether capability boundaries are already encoded in the hidden states of input questions.
Since hidden states are high-dimensional (e.g., 4096), limited samples may lead to sparsity issues that obscure statistical patterns (Vishwakarma, 2024; Köppen, 2000). To mitigate this, we first expand our dataset by including GSM8K (Cobbe et al., 2021) and Humanity’s Last Exam (HLE) (Phan et al., 2025). These two benchmarks provide complementary levels of difficulty (see Appendix A for full accuracy results): GSM8K is relatively easy, with an accuracy exceeding 90%; HLE is extremely difficult, with an accuracy below 10%. This contrast helps amplify the distinction between solvable and unsolvable questions, making capability boundaries more salient.
Next, we extract hidden states from the prefilling stage as it processes the input question (Yuan et al., 2024; Pope et al., 2023), using the hook method (Vig, 2019). Specifically, we focus on the hidden state of the last input token, which encapsulates the semantic content of the entire question (Radford et al., 2019). Finally, we apply simple linear classifiers to these hidden states, including Linear Discriminant Analysis (LDA) and Logistic Regression (LR) (Hastie, 2009). Using final answer correctness as labels, we test whether solvable and unsolvable questions can be linearly separated, thereby revealing the presence of capability boundaries in hidden states.
| LDA | LR | ||
| GPT-oss-20B | |||
| R1-Distill-Qwen3-8B | |||
| R1-Distill-Qwen-32B | |||
| QwQ-32B | |||
The left subfigure of Figure 5 visualizes the classification of solvable and unsolvable questions (see Appendix D for results of all LRMs), along with the capability boundary (blue line) determined by the decision boundary of linear classifier (e.g., LDA). We observe that the two classes of questions are clearly separated, under random train – test splits. To quantify this separation, Table 2 shows the classification accuracy, which is consistently close to and robust to the choice of classifier and hyperparameters (e.g., the regularization coefficient of LR). These results demonstrate that the features of capability boundaries are strongly encoded in the hidden states of the last input token.
Interestingly, we also observe that among solvable questions, some lie very close to the capability boundary, while others are far away. We compare these two cases by measuring their token usage during inference. Although LRMs are able to produce correct answers, the average token usage of the questions close to the boundary is – times longer than that far from the boundary (right subfigure of Figure 5). This further supports the validity of the identified capability boundary: solvable questions closer to the boundary require longer reasoning to reach the correct answer.
Therefore, the capability boundary is encoded in the hidden states of the last input token, and it also reflects the reasoning effort required to solve the question.
6 Reason with Capability Boundary
The analysis in the previous sections reveal that LRMs exhibit clear signs of self-awareness on capability boundaries: reasoning expressions expose confidence patterns that distinguish solvable from unsolvable problems, and hidden states encode boundary information even before reasoning begins. These findings suggest that capability boundaries are not only detectable but can also be harnessed to improve efficiency and reliability of LRMs’ response. In this section, we propose two strategies to reason with capability boundary and demonstrate their effectiveness.
6.1 Experimental Design
Strategies. Building on the evidence that self-awareness of capability boundaries emerges during test time, we propose two test-time monitoring strategies that leverage boundary-awareness to guide the reasoning process. The key idea is to monitor signals of capability boundaries on the fly, either from reasoning trajectories (black-box) or hidden states (white-box), and to adjust reasoning when questions are beyond the model’s capability. This approach allows LRMs to avoid unproductive reasoning and give more reliable response.
Reasoning Expression Monitoring (Monitor) is a black-box strategy that leverages expressions density during reasoning. The pipeline is as follows:
1.
Construct expression trajectories: continuously track the densities of confident and uncertain expressions (e.g., “I think that’s it” vs. “I might be wrong”) across reasoning steps to form expression density trajectories and .
2.
Detect unsolvable question: analyze these trajectories on the fly using quantitative indicators such as Confidence Differential (ConfDiff) and Confidence Curvature (ConfCurv) to determine whether a question lies within or beyond the capability boundary.
3.
Early stop unproductive reasoning and reprompt with suffix: if the trajectory indicates the question is beyond the capability boundary, we early terminate reasoning and reprompt the model to output a concise approach. This is implemented by appending a self-awareness prompt suffix (shown below) to the end of the question; otherwise, it continues with the standard reasoning process.
1.
Hidden States Monitoring (Monitor) is a white-box strategy that leverages the hidden states of the last input token. The pipeline is as follows:
1.
Extract hidden states: capture the hidden states of the last input token, which encodes the semantic representation of the entire question during the prefilling stage.
2.
Detect unsolvable question: apply simple linear classifiers (e.g., LDA or LR) trained to distinguish unsolvable from solvable questions. This is performed before reasoning begins.
3.
Avoid unproductive reasoning via constrained output prefix: if the classifier predicts that the question is beyond the capability boundary, the model abstains from full reasoning and instead outputs a concise possible approach. This is implemented by constraining the response to begin with a self-awareness output prefix (shown below); otherwise, the model proceeds with the standard reasoning process.
1.
Notably, both strategies preserve the model’s performance on solvable questions while providing more efficient and reliable responses on questions beyond its capability boundary.
Metrics. We adopt four metrics to systematically evaluate the two proposed strategies:
Accuracy (ACC) measures the proportion of correctly solved problems, serving as the standard indicator of reasoning quality (Xia et al., 2025).
Hard Abstention (HA) measures the model’s ability to recognize problems beyond its capability boundary. Different to abstention which focus on underspecification, outdated information, or noncompliance (Kirichenko et al., 2025; Madhusudhan et al., 2024; Brahman et al., 2024), HA specifically evaluates whether LRMs can explicitly abstain on unsolvable problems instead of producing incorrect answers.
Token usage measures the number of tokens consumed during reasoning and answering (Snell et al., 2024). This metric directly reflects the efficiency of the reasoning process.
Overflow measures the proportion of cases where the model’s output exceeds the context window, which identifies incomplete or abnormal responses (Ben-Kish et al., 2025).
Notably, ACC is evaluated on the full dataset, whereas the remaining three metrics are computed only on questions for which LRMs produce incorrect answers. For token usage and overflow, we vary the models’ context length to observe how our strategies scale in longer context scenarios.
Baseline. Prior work has explored improving reasoning performance by self-verification after completing the full Chain-of-Thought (Madhusudhan et al., 2024) or by monitoring undesirable behaviors during training (Baker et al., 2025). However, these approaches are orthogonal to our scenario: the former introduces additional computation after reasoning, while the latter does not address hard questions beyond the capability boundary.
Therefore, we adopt a baseline that improve efficiency by enabling the model to abstain. In particular, we consider BoostAbstention (Kirichenko et al., 2025) which crafts a system prompt that encourages the model to abstain in designated scenarios, showing effectiveness for both reasoning and standard LLMs. For our experiments, we modify the system prompt to better match the abstention scenarios of questions beyond capability boundary (see Appendix B for the full prompt).
6.2 Results
Table 3 compares the performance of the original models, the baseline BoostAbstention, and our proposed monitoring strategies. We highlight the following key observation.
| Context Length | 2048 | 4096 | ||||
| Metric | ACC (%) ↑ | HA (%) ↑ | Token ↓ | Overflow (%) ↓ | Token ↓ | Overflow (%) ↓ |
| GPT-oss-20B | ||||||
| +BoostAbstention | ||||||
| +Monitor | ||||||
| +Monitor | ||||||
| R1-Distill-Qwen3-8B | ||||||
| +BoostAbstention | ||||||
| +Monitor | ||||||
| +Monitor | ||||||
| R1-Distill-Qwen-32B | ||||||
| +BoostAbstention | ||||||
| +Monitor | ||||||
| +Monitor | ||||||
| QwQ-32B | ||||||
| +BoostAbstention | ||||||
| +Monitor | ||||||
| +Monitor | ||||||
Observation 1: our strategies enable LRMs to recognize unsolvable questions without compromising their accuracy on solvable ones.
Although the original models achieve acceptable Accuracy (ACC), their Hard Abstention (HA) is zero, indicating a complete lack of ability to abstain on questions beyond the capability boundary. Moreover, almost all Overflow is 100%, suggesting that most incorrect cases arise because LRMs struggling to solve a hard question until exceeding the context length.
Promisingly, both Monitor and Monitor achieve strong performance. First, their ACC remains nearly unchanged compared to the original models, showing that good reasoning performance is preserved. Meanwhile, the substantially high HA (almost 100%) demonstrates that LRMs can now effectively identify problems beyond the capability boundary.
Surprisingly, and contrary to the observation from Kirichenko et al. (2025), the baseline BoostAbstention has almost no effect in our scenario. This reveals that mathematical reasoning problems differ from outdated or unsafe requests: abstention cannot be induced simply by system prompts, but instead requires boundary signals encoded in hidden states or reasoning expressions.
Observation 2: our strategies enable LRMs to provide more efficient and reliable responses to unsolvable questions by offering concise solution approaches.
Once questions beyond the capability boundary are recognized, our strategies enable LRMs to improve efficiency and reliability by offering concise solution approaches. Figure 6 illustrates the responses produced by our strategies compared to the original models. The original models struggle with unsolvable questions and often fail to complete the reasoning process even after exhausting the context length. Such responses provide neither a correct answer nor valuable insights from the lengthy reasoning trace, while also wasting user time.
In contrast, our strategies detect these unsolvable questions and guide the model to explicitly acknowledge that they are beyond capability boundary, while offering a concise possible approach. This response paradigm provides users with possible hints for problem solving without wasting time on meaningless output. And importantly, it does not compromise the model’s ability to produce correct answers for solvable questions.
This explains why we observe large reductions in both token usage (up to 93.6%) and overflow (up to 100%) in Table 3, indicating that reasoning with capability boundaries greatly improves efficiency and mitigates incomplete outputs. Besides, this reduction becomes more evident with larger context length. For example, overflow reduces from 78.4% to 5.4% on GPT-oss-20B when the context length is extended from 2048 to 4096. This suggests that our strategies holds promise for optimizing long-context reasoning tasks (Ling et al., 2025; Kuratov et al., 2024).
Observation 3: LRMs vary in reasoning behavior correction, with some good at instruction following and others more prone to output guidance.
We compare the performance of Monitor and Monitor across different LRMs. For GPT-oss-20B, Monitor achieves lower token usage than Monitor. We find this model is resistant to output guidance via constrained output prefix: although it sometimes begins by generating a solution approach, it tends to later negate this and restart the reasoning process, resulting in long and unproductive outputs. In contrast, its instruction-following ability is relatively strong, and thus reprompting with the self-awareness prompt suffix yields lower token usage.
Conversely, the other three LRMs show weaker ability of instruction following but more prone to output guidance, making Monitor more effective than Monitor. This suggest that, in addition to the black-box versus white-box distinction, the choice of optimization strategy in practice should also consider the behavioral tendencies of the specific LRM.
7 Conclusion
In this work, we investigate whether Large Reasoning Models (LRMs) possess self-awareness of their capability boundaries. Through both black-box and white-box analyses, we found consistent evidence that LRMs can distinguish unsolvable from solvable questions: reasoning expressions reveal separable expression trajectories, while hidden states of the last input token encode boundary information even before reasoning begins. Building on these insights, we proposed two simple yet effective test-time monitoring strategies that leverage boundary signals to prevent unproductive reasoning for unsolvable questions. Our experiments show that these strategies preserve accuracy on solvable problems while substantially improving efficiency and reliability on unsolvable ones.
Appendix A Datasets, Models, and Performance
In this section, we provide the specific description of models and datasets evaluated in our work. And we provide the performance of each model on each dataset.
We evaluate the following LRMs:
GPT-oss-20B (Agarwal et al., 2025): An open-source large reasoning model with 20B parameters, designed as a general-purpose backbone for multi-step reasoning tasks.
DeepSeek-R1-0528-Qwen3-8B (Guo et al., 2025): A distilled variant of DeepSeek-R1 based on Qwen3-8B, post-trained with chain-of-thought supervision. It achieves state-of-the-art performance among open-source models on mathematical reasoning benchmarks.
DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025): A 32B-parameter distilled model obtained by transferring reasoning abilities from DeepSeek-R1 into the Qwen2.5-32B backbone, yielding strong performance across multiple benchmarks.
QwQ-32B (Qwen Team, 2024): A medium-sized reasoning model from the Qwen series, competitive with other state-of-the-art reasoning-focused models on math and logical reasoning tasks.
We evaluate the following Math datasets:
AIME’25 (MAA, 2025): The 2025 American Invitational Mathematics Examination. It consists of 30 integer-answer questions with increasing difficulty, used as a challenging benchmark for mathematical reasoning.
AIME’24 (MAA, 2024): The 2024 American Invitational Mathematics Examination. Similar in format to AIME’25, it consists of 30 problems and provides a prior-year benchmark for evaluating model robustness across exam editions.
HMMT (HMMT, 2025): The Harvard–MIT Mathematics Tournament (February 2025 session). The dataset contains 30 questions sourced from the original competition, which were extracted, converted to LaTeX, and verified. These problems feature complex multi-step reasoning across various mathematical domains.
AMC’23 (MAA, 2023): The 2023 American Mathematics Competition (AMC 10A). It consists of 30 questions, providing a lighter but diverse benchmark compared to AIME and HMMT.
GSM8K (Cobbe et al., 2021): A large-scale dataset of grade-school math word problems requiring multi-step arithmetic reasoning. We evaluate on approximately 1.5k problems from the test set. It is widely used to evaluate the basic reasoning competence of LRMs, with accuracy typically exceeding 90%.
Humanity’s Last Exam (HLE) (Phan et al., 2025): A multi-modal benchmark at the frontier of human knowledge, designed with broad subject coverage across mathematics, humanities, and natural sciences. We use approximately 1.5k text-only questions (excluding image-based problems) from this dataset. Model accuracy on this dataset is typically below 10%, making it a strong stress test for reasoning beyond capability boundaries.
The performance of LRMs on these Math benchmarks are shown in the following Table 4:
| AIME’25 | AIME’24 | AMC’23 | HMMT | GSM8K | HLE | |
| GPT-oss-20B | ||||||
| R1-Distill-Qwen3-8B | ||||||
| R1-Distill-Qwen-32B | ||||||
| QwQ-32B |
Appendix B System Prompt of BoostAbstention
We consider BoostAbstention (Kirichenko et al., 2025) which crafts a system prompt that encourages the model to abstain in designated scenarios, showing effectiveness for both reasoning and standard LLMs. For our experiments, we modify the system prompt to better match the abstention scenarios of questions beyond capability boundary:
Appendix C Confident and Uncertain Expressions
Figure 7 shows the confident and uncertain expression that frequently appear in the reasoning process. The expressions are summarized manually and checked by LLM. In the experiment, when matching these expressions, the Uncertain Expression is matched first, followed by the Confident Expression. Additionally, both the position of the first token of an expression and the token length of the entire expression are considered.
Appendix D Capability Boundary Revealed by Hidden States
Figure 8 visualizes the classification of solvable and unsolvable questions of all LRMs, along with the capability boundary (blue line) determined by the decision boundary of linear classifier (e.g., LDA). We observe that the two classes of questions are clearly separated, under random train–test splits. These results demonstrate that the features of capability boundaries are strongly encoded in the hidden states of the last input token, and is valid across different LRMs.
Appendix E Unproductive Reasoning
Figure 9 and Figure 10, demonstrate representative unproductive reasoning phenomena in large language models, including repetitive looping and error accumulation, which impair their capacity for sound and consistent reasoning in complex problem-solving scenarios. They respectively show how models get stuck in repetitive deductions (Figure 9) and how errors compound during formula application and result derivation (Figure 10). In mathematical problem reasoning, the former presents repeated deduction of the same uncertain conclusion, while the latter sees continuous error accumulation due to flawed initial methods.