### v3

# Stop Before You Fail: Operational Capability Boundaries for Mitigating Unproductive Reasoning in Large Reasoning Models

Abstract
Current answering paradigms for Large Reasoning Models (LRMs) often fail to account for the fact that some questions may lie beyond the model’s operational capability boundary, leading to long but unproductive reasoning. In this paper, we study whether LRMs expose early signals predictive of such cases, and whether these signals can be used to mitigate unproductive reasoning. In black-box settings, we find that reasoning expressions contain failure-predictive signals. In white-box settings, we show that the hidden states of the last input token contain information that is predictive of whether a question will not be solved correctly under our evaluation setup. Building on these observations, we propose two test-time monitoring strategies: reasoning expression monitoring and hidden states monitoring, that reduce token usage by 62.7–93.6%, substantially improving efficiency and reliability while largely preserving accuracy.111We open-source our code.
Stop Before You Fail: Operational Capability Boundaries for Mitigating Unproductive Reasoning in Large Reasoning Models
Qingjie Zhang1, Yujia Fu1, Yang Wang2, Liu Yan2,
Tao Wei2, Ke Xu1, Minlie Huang1, and Han Qiu1††thanks: The corresponding author
1Tsinghua University, 2Ant Group
Emails:{qj-zhang24@mails., qiuhan@}tsinghua.edu.cn
1 Introduction
Large Reasoning Models (LRMs) have demonstrated remarkable capabilities on complex reasoning tasks such as mathematics (Guo et al., 2025; Jaech et al., 2024; Ahn et al., 2024). However, when confronted with hard questions, they frequently exhibit unproductive reasoning, manifested as repetitive looping or error accumulation until context limits are exhausted (Zhou et al., 2025; Chen et al., 2025b; Yao et al., 2025; Mukherjee et al., 2025) (as shown in Appendix K).
This phenomenon suggests an important limitation of current answering paradigms: they often fail to account for the fact that some questions are likely to fall beyond the model’s operational capability boundary under a fixed evaluation setup, leading to long yet ultimately unproductive reasoning traces. If such cases can be identified early, reasoning could become more reliable and efficient: instead of continuing low-value deliberation, the model may abstain from full reasoning and provide a concise partial approach.
Existing approaches to studying reasoning limits have primarily taken an external perspective, for example by constructing increasingly challenging benchmarks such as Humanity’s Last Exam (HLE) (Phan et al., 2025), or by designing real-time routers that estimate question difficulty and route harder questions to more capable models (OpenAI, 2025). In contrast, relatively little work has examined whether such limits are reflected in the model’s internal signals, despite the importance of this question for both understanding reasoning behavior and further improving reliability and efficiency at test time.
Accordingly, this paper studies whether LRMs expose early signals predictive of eventual failure, and whether such signals can be used to mitigate unproductive reasoning. We use the term capability boundary in an operational sense: under a fixed evaluation setup, it refers to the empirical separation between questions the model eventually solves and those it does not. Figure 1 provides an overview of our approach. From both black-box and white-box perspectives, we identify signals that help distinguish unsolvable from solvable questions. Leveraging these signals, we enable LRMs to assess, on the fly, whether a question is likely to remain unsolved under the current setup, thereby improving the efficiency and reliability of reasoning. We summarize our contributions as:
We provide empirical evidence that LRMs’ reasoning expressions are associated with the correctness of final answers, revealing language patterns that are predictive of eventual failure.
From a black-box perspective, we characterize failure-predictive reasoning trajectories. We identify a diverging pattern: an accelerated growing confidence trajectory for solved questions, but a convergent uncertainty trajectory for unsolved questions.
From a white-box perspective, we show that the hidden states of the last input token contain information predictive of whether a question will eventually be solved correctly. In particular, solvable and unsolvable questions are close to linearly separable in the hidden space, even before reasoning begins.
We design two test-time monitoring strategies: reasoning expression monitoring (black-box) and hidden states monitoring (white-box). Experiments show that these strategies substantially improve efficiency and reliability while largely preserving accuracy, reducing token usage by –.
2 Related Work
From knowledge boundary to capability boundary. Recent studies have examined the knowledge boundaries of Large Language Models (LLMs), focusing on the limits of factual knowledge encoded in parameters (Yin et al., 2024; Li et al., 2024). For instance, Wen et al. (2024) investigate LLMs’ perception of knowledge boundary through semi-open-ended question answering. Deng et al. (2025) unveil knowledge boundary of LLMs for trustworthy information access. These works primarily emphasize factual coverage, with retrieval-augmented generation (RAG) being a common solution to extend the accessible knowledge space (Gao et al., 2023; Lewis et al., 2020).
By contrast, the capability boundary concerns what reasoning tasks LLMs can reliably solve (Chen et al., 2025a; Xue et al., 2025; Chen et al., 2024a). This notion is broader yet rarely formalized. Existing efforts typically approximate it externally: either by constructing more challenging benchmarks such as Humanity’s Last Exam (HLE) (Phan et al., 2025), or by designing routing mechanisms such as GPT-5 router that delegate harder questions to more capable models (OpenAI, 2025). In contrast, we examine whether capability boundaries are reflected in models’ internal signals.
Interpretability of reasoning expressions and hidden states. Interpretability aims to uncover how LLMs understand and execute tasks, providing insights that enhance their reliability and efficiency. Recent studies have approached from both reasoning expressions and hidden states. On the reasoning expressions side, Guo et al. (2025) highlight the anthropomorphic cues, “aha moments”, as a key role in reasoning. Qian et al. (2025) tracks mutual information along the generation and finds sharp surges at specific thinking tokens (e.g., “Hmm”, “Wait”, “Therefore”). Bogdan et al. (2025) frame sentence-level importance as a counterfactual influence problem, showing that only a few critical utterances steer the reasoning trajectory.
On the hidden states side, Wendler et al. (2024) probe multilingual Llama models and reveal English as the pivot language in the internal structure. Fan et al. (2024) demonstrate that certain intermediate layers already encode sufficient predictive features, implying redundancy in the full forward pass. Zhang et al. (2025) train linear probes on hidden states to verify intermediate answers, anticipating future correctness. Building on both perspectives, our work analyzes reasoning expressions and hidden states to study signals related to the operational capability boundary.
Efficient reasoning: from solvable questions to unsolvable questions. A range of recent works examine the phenomenon of overthinking (Sui et al., 2025; Hou et al., 2025). Chen et al. (2024b) show that LRMs often allocate excessive computation to easy questions, while Zhang et al. (2024) report that multi-turn self-correction can lead to unnecessary cost. To mitigate such inefficiency, Han et al. (2025) compresses unnecessary reasoning by including a reasonable token budget in the prompt. Wang et al. (2025) demonstrate that suppressing frequent reflective tokens such as “Wait”, “Hmm”, or “Alternatively” improves reasoning efficiency without harming accuracy. Yang et al. (2025a); Fu et al. (2025, 2024) introduce mechanisms to decide during reasoning whether further steps are unnecessary, allowing the model to terminate early and output the final answer directly.
However, these works focus on cases where LRMs can solve a question but still waste computation by reasoning for too long. By contrast, our work addresses a complementary and less-studied scenario: questions that are likely to fall beyond the model’s operational capability boundary, where LRMs may nevertheless continue reasoning unproductively despite being unlikely to arrive at a correct answer under the current setup.
3 LRMs Expose Signals of Failure
We begin our exploration by evaluating GPT-oss-20B (Agarwal et al., 2025), DeepSeek-R1-0528-Qwen3-8B (abbreviated as R1-Distill-Qwen3-8B) (Guo et al., 2025), DeepSeek-R1-Distill-Qwen-32B (abbreviated as R1-Distill-Qwen-32B) (Guo et al., 2025), and QwQ-32B (Qwen Team, 2024) on several standard mathematical benchmarks, including AIME’25 (MAA, 2025), AIME’24 (MAA, 2024), HMMT (HMMT, 2025), and AMC’23 (MAA, 2023) (see Appendix A for detailed descriptions of datasets and models).
Inspired by works which monitors LRMs’ reasoning (Baker et al., 2025; He et al., 2025; Luo et al., 2024), we examine the reasoning processes (e.g., the segments wrapped within <think> in DeepSeek-like models) to look for internalized assessments regarding whether the input question can be correctly solved. As illustrated in Figure 2, we observe that such signals, expressed as anthropomorphic tones (Guo et al., 2025; Yang et al., 2025b), emerge throughout the reasoning process, and become especially evident in its final stages. Toward the end of reasoning, models often produce explicit statements that reflect their confidence level. Strikingly, these expressed beliefs tend to align with the correctness of their final answers. For example, when the reasoning contains more confident expressions such as “So, no mistake” or “I think it’s correct”, the final answers are typically correct. Conversely, when uncertain expressions like “My initial assumption is wrong” or “I’m not 100% sure” appear, the final answers are incorrect.
This observation motivates us to quantify the alignment between expressed belief and answer correctness, which we formalize through:
| Can % | (1) | |||
| Cannot % |
Specifically, Can % is the proportion of questions for which the model expresses more confidence during reasoning and indeed produces a correct answer, while Cannot % is the proportion of questions for which the model expresses more uncertainty and indeed fails to provide the correct answer.
| Can % | Cannot % | |
| GPT-oss-20B | ||
| R1-Distill-Qwen3-8B | ||
| R1-Distill-Qwen-32B | ||
| QwQ-32B |
Table 1 shows the results. We observe that LRMs’ expressed beliefs align with the correctness of their final answers. In particular, for questions where LRMs express more uncertainty during reasoning, the final answers are consistently incorrect (Cannot % = 100). This phenomenon is observed across multiple models, suggesting that LRMs expose signals of failure.
4 Capability Boundary in Expressions
Motivated by the observations in the previous section, we hypothesize that LRMs’ reasoning expressions may contain signals related to the operational capability boundary. We therefore systematically investigate the dynamics of confident and uncertain expressions throughout the reasoning process.
To ensure that the analysis is not tied to any particular LRM’s reasoning style, we first extract confident and uncertain expressions separately for each model, based on its characteristic reasoning patterns (see the detailed extraction pipeline in Appendix C). Such anthropomorphic expressions are commonly observed in LRM reasoning traces and have been discussed in prior work as informative features of the reasoning process (Yang et al., 2025b; Guo et al., 2025; Qian et al., 2025).
Then, we uniformly divide the reasoning process into multiple stages (e.g., 50 stages) and compute the density of confident and uncertain expressions at each stage. Finally, concatenating these values yields the expression density trajectories, denoted as for confident expressions and for uncertain expressions. These trajectories capture how the density of each type of expression evolves over the course of reasoning.
We plot and in Figure 4, separately for cases with correct and wrong final answers. We observe that LRMs exhibit clear differences in their reasoning expression trajectories depending on final answer correctness. When the final answer is correct, the confident trajectory lies above the uncertain trajectory and shows an accelerating increasing trend. In contrast, when the final answer is incorrect, the uncertain trajectory dominates over the confident trajectory and tends to converge. This contrast suggests that reasoning expression trajectories contain useful signals (blue line in Figure 4) for distinguishing questions that are eventually solved from those that remain unsolved under the current setup. In particular, the dominant trajectory tends to be more concave for solvable questions and more convex for unsolvable ones.
To formalize these patterns, we design two indicators for determining whether a question is likely to lie within or beyond the capability boundary.
Confidence Differential (ConfDiff) measures the sign of gap between uncertain and confident trajectory along the reasoning process. Specifically, we accumulate the sign up to a given reasoning stage percentage , and treat the question as beyond the capability boundary if the cumulative value exceeds a threshold :
Confidence Curvature (ConfCurv) computes the second derivative of the expression density trajectory to capture its convexity (Boyd and Vandenberghe, 2004). A concave trajectory corresponds to questions within the capability boundary, while a convex trajectory indicates questions beyond it. Also, we accumulate the sign of convexity up to a reasoning stage percentage and treat the question as beyond the capability boundary if the value exceeds a threshold :
Figure 3 shows that the two proposed indicators can reliably distinguish unsolvable from solvable questions, achieving high accuracy across different LRMs throughout the reasoning stage. Between the two indicators, ConfDiff is more reliable as it achieves higher accuracy and exhibits smaller fluctuations during the reasoning process. Moreover, strong predictive accuracy emerges early, e.g., as early as 2% for GPT-oss-20B. Interestingly, the gradually declining yellow curve of this model suggests that these signals may be more salient at the beginning than at the end of reasoning. One possible explanation is that repetitive looping and error accumulation in later stages make the trajectories noisier (Yao et al., 2025; Mukherjee et al., 2025). Overall, these results show that reasoning expressions contain early signals related to the operational capability boundary and predictive of eventual failure.
5 Capability Boundary in Hidden States
Following the observation that reasoning expressions expose early signals related to the operational capability boundary, it is natural to ask whether similar signals may emerge even earlier, potentially before the model begins reasoning. We therefore investigate whether hidden states of the input question already contain information predictive of eventual success or failure.
| LDA | LR | ||
| GPT-oss-20B | |||
| R1-Distill-Qwen3-8B | |||
| R1-Distill-Qwen-32B | |||
| QwQ-32B | |||
Since hidden states are high-dimensional (e.g., 4096), limited samples may lead to sparsity issues that obscure statistical patterns (Vishwakarma, 2024; Köppen, 2000). We mix the mathmatical benchmarks used in the previous sections and randomly split it into an 80% training set and a 20% test set, with balanced labels for eventually unsolved and solved questions.
Next, we extract hidden states from the prefilling stage as it processes the input question (Yuan et al., 2024; Pope et al., 2023), using the hook method (Vig, 2019). Specifically, we focus on the hidden state of the last input token, which encapsulates the semantic content of the entire question (Dong et al., 2025; Radford et al., 2019). Finally, we apply simple linear classifiers to these hidden states, including Linear Discriminant Analysis (LDA) and Logistic Regression (LR) (Hastie, 2009). Using final answer correctness as labels, we test whether eventually solved and unsolved questions can be linearly separated, i.e., whether last token hidden states contain information predictive of eventual correctness under our setup.
The left subfigure of Figure 5 visualizes the classification of solvable and unsolvable questions (see Appendix H for results of all LRMs), together with the decision boundary (blue line) of a linear classifier (e.g., LDA). We observe that the two classes of questions are clearly separated under random train–test splits. To quantify this separation, Table 2 reports the classification accuracy, which is consistently close to and robust to the choice of classifier and hyperparameters (e.g., the regularization coefficient of LR). These results suggest that the hidden states of the last input token contain information that is predictive of eventual correctness, and thus closely related to the operational capability boundary under our evaluation setup.
At the same time, these signals should be interpreted with care: they may partially reflect question difficulty rather than a purely intrinsic notion of capability boundary. We provide additional analyses and discussion of this issue in Appendix E.
Interestingly, we also observe that among eventually solved questions, some lie very close to the classifier decision boundary, while others lie farther away. We quantify this distance by the probability difference assigned by the LDA classifier to the solved and unsolved classes (Hwang et al., 2023). We then compare these groups in terms of token usage during inference. Although LRMs answer both groups correctly, questions closer to the boundary require on average – more tokens than those farther from the boundary (right subfigure of Figure 5). This suggests that the hidden state signals are related not only to eventual correctness, but also to the amount of reasoning effort required to obtain a correct answer.
Overall, these results indicate that the hidden states of the last input token contain useful signals related to the operational capability boundary, and that these signals also correlate with the reasoning effort required to solve a question.
Moreover, because these signals are extracted from reasoning expressions and hidden states rather than task-specific heuristics, they may extend beyond mathematical reasoning. We provide an analysis on a coding task in Appendix G.
6 Reason with Capability Boundary
Because signals related to the operational capability boundary emerge early, they can be used to improve the efficiency and reliability of LRMs’ responses. In this section, we propose two test-time monitoring strategies that leverage such signals and demonstrate their effectiveness.
6.1 Experimental Design
Strategies. The key idea is to monitor signals of capability boundaries on the fly, either from reasoning trajectories (black-box) or hidden states (white-box), and to adjust reasoning when a question lies beyond the model’s effective solving regime.
Reason Expression Monitoring (Monitor) is a black-box strategy that leverages expressions density during reasoning. The pipeline is:
1. Construct expression trajectories: continuously track the densities of confident and uncertain expressions (e.g., “I think that’s it” vs. “I might be wrong”) across reasoning steps to form expression density trajectories and .
2. Predict likely-unsolved question: analyze these expression trajectories on the fly using quantitative indicators such as Confidence Differential (ConfDiff) to determine whether a question is likely to remain unsolved.
3. Early stop unproductive reasoning and reprompt with suffix: if the trajectory indicates the question is likely to remain unsolved, we early terminate reasoning and reprompt the model to output a concise approach. This is done by appending a prompt suffix at the end of the question:
Hidden States Monitoring (Monitor) is a white-box strategy that leverages the hidden states of the last input token. The pipeline is as follows:
1. Extract hidden states: capture the hidden states of the last input token, which encodes the semantic representation of the entire question during the prefilling stage.
2. Predict likely-unsolved question: apply simple linear classifiers (e.g., LDA or LR) trained to distinguish eventually unsolved from solved questions. This prediction is made before reasoning begins.
3. Avoid unproductive reasoning via constrained output prefix: if the classifier predicts that the question is likely to remain unsolved under the current setup, the model abstains from full reasoning and instead outputs a concise possible approach. This is implemented by constraining the response to begin with an output prefix (see Appendix F for the ablation study):
| Context Length | 2048 | 4096 | ||||
| Metric | ACC (%) ↑ | HA (%) ↑ | Token ↓ | Overflow (%) ↓ | Token ↓ | Overflow (%) ↓ |
| GPT-oss-20B | ||||||
| +BoostAbstention | ||||||
| +Monitor | ||||||
| +Monitor | ||||||
| R1-Distill-Qwen3-8B | ||||||
| +BoostAbstention | ||||||
| +Monitor | ||||||
| +Monitor | ||||||
| R1-Distill-Qwen-32B | ||||||
| +BoostAbstention | ||||||
| +Monitor | ||||||
| +Monitor | ||||||
| QwQ-32B | ||||||
| +BoostAbstention | ||||||
| +Monitor | ||||||
| +Monitor | ||||||
Metrics. We adopt four metrics to systematically evaluate the two proposed strategies:
Accuracy (ACC) measures the proportion of correct answers, serving as the standard indicator of reasoning quality (Xia et al., 2025).
Hard Abstention (HA) measures the model’s ability to recognize questions beyond its capability boundary. Different to abstention which focus on underspecification, outdated information, or noncompliance (Kirichenko et al., 2025; Madhusudhan et al., 2024; Brahman et al., 2024), HA specifically evaluates whether LRMs can explicitly abstain on unsolvable questions instead of producing incorrect answers.
Token usage measures the number of tokens consumed during reasoning and answering (Snell et al., 2024). This metric directly reflects the efficiency of the reasoning process.
Overflow measures the proportion of cases where the model’s output exceeds the context window, which identifies incomplete or abnormal responses (Ben-Kish et al., 2025).
Notably, ACC is evaluated on the full dataset, whereas the remaining three metrics are computed only on questions for which LRMs produce incorrect answers. For token usage and overflow, we vary the models’ context length to observe how our strategies scale in longer context scenarios.
Baseline. Prior work has explored improving performance by self-verification after generating the answer (Madhusudhan et al., 2024; Kadavath et al., 2022) or by monitoring uncertainty and other undesirable behaviors (Baker et al., 2025; Farquhar et al., 2024; Kuhn et al., 2023). However, these approaches are orthogonal to our scenario: the former introduces additional computation after producing an answer, while the latter does not address eventually unsolved questions (see Appendix B).
Therefore, we adopt a baseline that improve efficiency by enabling the model to abstain. In particular, we consider BoostAbstention (Kirichenko et al., 2025) which crafts a system prompt that encourages the model to abstain in designated scenarios, showing effectiveness for both reasoning and standard LLMs. For our experiments, we modify the system prompt to better match the abstention scenarios of eventually unsolved questions (see Appendix D for the full prompt).
6.2 Results
Table 3 compares the performance of the original models, the baseline BoostAbstention, and our proposed monitoring strategies. We highlight the following key observation.
Observation 1: our strategies enable LRMs to recognize eventually unsolved questions while maintaining comparable accuracy on solved ones. Although the original models achieve acceptable Accuracy (ACC), their Hard Abstention (HA) is zero, indicating a complete lack of ability to abstain on questions beyond the capability boundary. Moreover, almost all Overflow is 100%, suggesting that most incorrect cases arise because LRMs struggling to solve a hard question until context limit.
Promisingly, both strategies achieve strong performance. First, their ACC remains nearly unchanged compared to the original models, showing that good reasoning performance is preserved. Meanwhile, the substantially high HA (almost 100%) demonstrates that LRMs can now effectively identify likely-unsolved questions.
Surprisingly, and contrary to the observation from Kirichenko et al. (2025), the baseline BoostAbstention has almost no effect in our scenario. This reveals that mathematical reasoning problems differ from outdated or unsafe requests: abstention cannot be induced simply by system prompts, but instead requires boundary signals encoded in hidden states or reasoning expressions.
Observation 2: our strategies enable LRMs to provide more efficient and reliable responses to eventually unsolved questions by offering concise solution approaches. Figure 6 illustrates the responses produced by our strategies compared to the original models. The original models struggle with eventually unsolved questions and fail to complete the reasoning process even after exhausting the context length. Such responses provide neither a correct answer nor valuable insights from the lengthy reasoning trace, while also wasting time.
In contrast, our strategies detect these unsolved questions and guide the model to explicitly acknowledge that they are beyond capability boundary, while offering a concise possible approach. This response paradigm provides users with possible hints for problem solving without wasting time on meaningless output. And importantly, it does not compromise the model’s ability to produce correct answers for solvable questions.
This explains why we observe large reductions in both token usage (up to 93.6%) and overflow (up to 100%) in Table 3, indicating that reasoning with capability boundaries greatly improves efficiency and mitigates incomplete outputs. Besides, this reduction becomes more evident with larger context length. For example, overflow reduces from 78.4% to 5.4% on GPT-oss-20B when the context length is extended from 2048 to 4096. This suggests that our strategies hold promise for optimizing long-context reasoning tasks (Ling et al., 2025; Kuratov et al., 2024). We provide in Appendix I an additional analysis with the context length set to maximum.
Observation 3: LRMs vary in reasoning behavior correction, with some good at instruction following and others more prone to output guidance. We compare the performance of Monitor and Monitor across different LRMs. For GPT-oss-20B, Monitor achieves lower token usage than Monitor. We find this model is resistant to output guidance via constrained output prefix: although it sometimes begins by generating a solution approach, it tends to later negate this and restart the reasoning process, resulting in long and unproductive outputs. In contrast, its instruction-following ability is relatively strong, and thus reprompting with the prompt suffix via Monitor yields lower token usage.
Conversely, the other three LRMs show weaker ability of instruction following but more prone to output guidance, making Monitor more effective than Monitor. This suggest that, in addition to the black-box versus white-box distinction, the choice of optimization strategy in practice should also consider LRMs’ behavioral tendencies.
7 Conclusion
We study whether signals related to the operational capability boundary can help mitigate unproductive reasoning in LRMs. Across both black-box and white-box analyses, we find that LRMs expose such signals early: reasoning expressions differ systematically between eventually solved and unsolved questions, while the hidden states of the last input token are predictive of eventual correctness even before reasoning begins. Based on these findings, we propose two test-time monitoring strategies that reduce unproductive reasoning. Experiments show that they largely preserve accuracy on solvable questions while substantially improving efficiency and reliability on questions lie beyond the operational capability boundary.
Limitations
Interpretation of capability boundaries. In this paper, we use capability boundary in an operational sense, namely as the empirical separation between questions a model eventually solves and those it does not under a fixed evaluation setup. Accordingly, the signals we identify should be interpreted as predictive correlates of eventual failure rather than definitive evidence of a formally defined intrinsic capability boundary. In particular, these signals may partially reflect question difficulty. Although we attempt to mitigate this concern through per-benchmark and cross-benchmark analyses (Appendix E), the extent to which these signals capture a general notion of capability boundary remains an open question. Developing a more principled formalization of capability boundaries (e.g., disentangling question difficulty) is an important direction for future work.
Scope of tasks and evaluation domains. Our evaluation focuses primarily on mathematical benchmarks, which allow systematic analysis of unsolvability signals but may not fully represent other reasoning scenarios. While our experiments on coding tasks suggest similar results (Appendix G), the generality of capability boundary signals across broader domains (e.g., complex decision making or scientific reasoning) remains to be established. Future work should investigate whether similar strategies apply across more diverse tasks.
Ethics Statement
ACL Ethics Policy is respected in this work. It studies LRMs’ capability boundaries, aiming to improve efficiency and reliability. No human subjects, sensitive personal data, or harmful content were involved. The proposed methods are intended for research purposes only and do not pose foreseeable risks of misuse.
Appendix A Datasets and Models
In this section, we provide the specific description of models and datasets evaluated in our work.
We evaluate the following LRMs:
GPT-oss-20B (Agarwal et al., 2025): An open-source large reasoning model with 20B parameters, designed as a general-purpose backbone for multi-step reasoning tasks.
DeepSeek-R1-0528-Qwen3-8B (Guo et al., 2025): A distilled variant of DeepSeek-R1 based on Qwen3-8B, post-trained with chain-of-thought supervision. It achieves state-of-the-art performance among open-source models on mathematical reasoning benchmarks.
DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025): A 32B-parameter distilled model obtained by transferring reasoning abilities from DeepSeek-R1 into the Qwen2.5-32B backbone, yielding strong performance across benchmarks.
QwQ-32B (Qwen Team, 2024): A medium-sized reasoning model from the Qwen series, competitive with other state-of-the-art reasoning-focused models on math and logical reasoning tasks.
For the sampling settings, we use greedy decoding to ensure reproducibility of results. For the maximum sequence length, we follow the specifications in the HuggingFace model cards: 128k for GPT-oss-20B, 64k for DeepSeek-R1-0528-Qwen3-8B, and 32k for both DeepSeek-R1-Distill-Qwen-32B and QwQ-32B.
We evaluate the following Math datasets:
AIME’25 (MAA, 2025): The 2025 American Invitational Mathematics Examination. It consists of 30 integer-answer questions with increasing difficulty, used as a challenging benchmark for mathematical reasoning.
AIME’24 (MAA, 2024): The 2024 American Invitational Mathematics Examination. Similar in format to AIME’25, it consists of 30 problems and provides a prior-year benchmark for evaluating model robustness across editions.
HMMT (HMMT, 2025): The Harvard–MIT Mathematics Tournament (February 2025 session). The dataset contains 30 questions sourced from the original competition, which were extracted, converted to LaTeX, and verified. These problems feature complex multi-step reasoning across various mathematical domains.
AMC’23 (MAA, 2023): The 2023 American Mathematics Competition (AMC 10A). The dataset consists of 30 questions, providing a lighter but diverse benchmark compared to AIME and HMMT.
GSM8K (Cobbe et al., 2021): A large-scale dataset of grade-school math word problems requiring multi-step arithmetic reasoning. We evaluate on approximately 1.5k problems from the test set. It is widely used to evaluate the basic reasoning competence of LRMs.
Humanity’s Last Exam (HLE) (Phan et al., 2025): A multi-modal benchmark at the frontier of human knowledge, designed with broad subject coverage across mathematics, humanities, and natural sciences. We use approximately 1.5k text-only questions (excluding image-based problems) from this dataset.
Appendix B Limitations of Some Prior Work
Some prior works evaluate model confidence through self-verification (Kadavath et al., 2022) or uncertainty estimation (Farquhar et al., 2024; Kuhn et al., 2023). In this section, we show that these approaches are less effective than our proposed monitoring strategies for identifying unsolvable questions with comparable accuracy on solvable ones. We first clarify how we adapt these methods to fit our task.
Self-verification (SV). For each question, we ask the model to evaluate whether its own predicted answer is correct using the prompt template:
This follows prior work (Kadavath et al., 2022) which study whether language models can evaluate the validity of their own claims by asking models to first propose answers, and then to evaluate that their answers are correct. If the model judges its own incorrect answer to be incorrect, we count this as a hard abstention (HA). Conversely, if the model judges its correct answer to be incorrect, this reduces the overall task accuracy (Acc). Formally,
| Model judges the answer to be correct | (4) | |||
| Model judges the answer to be incorrect | (5) | |||
Uncertainty-estimation (UE). For each question, we compute the semantic entropy of the model’s generated answer, where a higher entropy indicates greater uncertainty. This follows prior work (Farquhar et al., 2024; Kuhn et al., 2023), which introduces semantic entropy as an uncertainty measure that captures linguistic invariances induced by shared meanings. When the entropy exceeds a certain threshold, the model is considered uncertain about its answer. Similar as before, we define:
| Method | ACC (%) ↑ | HA (%) ↑ |
| GPT-oss-20B | 74.6 | 0 |
| +SV | 32.1 | 46.0 |
| +UE | 50.0 | 29.0 |
| +Monitor | 74.6 | 100 |
| +Monitor | 73.7 | 98.9 |
Results. We provide the results of these two methods alongside our Monitor and Monitor in Table 4. Although both methods provide models with a certain degree of hard abstention capability, they also incorrectly classify many solvable questions as unsolvable, which significantly reduces overall accuracy. In contrast, our approaches enable LRMs to reliably identify unsolvable questions without compromising accuracy on solvable ones.
Appendix C Confident and Uncertain Expressions
To ensure that our analysis is not tied to particular LRM’s reasoning style, we extract confident and uncertain expressions individually for each model, based on its own characteristic reasoning patterns. The extraction pipeline is outlined below.
Step 1: Identify all anthropomorphic expressions from the model’s reasoning trace. Prior work (Guo et al., 2025; Qian et al., 2025) show that the anthropomorphic expressions (e.g., “wait”) are crucial for LRM reasoning performance, in contrast to purely neutral mathematical steps (e.g., “11 + 5 = 16”). Following this insight, we curate a set of anthropomorphic expressions that the LRM commonly uses. And we further verify with GPT-5 that these expressions indeed convey anthropomorphic tone rather than purely propositional content.
Step 2: Select expressions whose occurrence frequency is most predictive of answer correctness. For each question, we count the occurrences of all extracted expressions within reasoning. Using the final answer correctness as the label and the occurrence counts for each expressions as features, we apply feature-importance analysis methods (Wang et al., 2024), including SHAP and Permutation Importance (PI), to identify the subset of expressions most informative for predicting whether the model’s final answer is correct. As shown in Table 5 (an example for GPT-oss-20B), we select expressions such as “wait”, “yes”, and “exactly” which rank among the most predictive, whereas others (e.g., “so”) exhibit negligible predictive value and are therefore excluded.
| Expression | SHAP | PI | Total |
| wait | 1.00 | 1.00 | 2.00 |
| yes | 0.33 | 0.73 | 1.06 |
| exactly | 0.26 | 0.73 | 0.99 |
| so | 0.01 | 0.00 | 0.01 |
Step 3: Divide the selected expressions into confidence and uncertainty categories. This step is grounded in the connection between epistemic markers and expressed uncertainty. Prior work (Liu et al., 2025; Yona et al., 2024) show that epistemic markers reflect LLMs’ internal confidence, analogous to how humans use expressions such as “I am not sure” or “it is unlikely that” to convey uncertainty. Following this insight, we classify each selected expression as reflecting confidence or uncertainty (further validated by GPT-5).
Figure 7 shows the full list of expressions of all LRMs. When matching these expressions, the Uncertain Expression is matched first, followed by the Confident Expression. Additionally, both the position of the first token of an expression and the token length of the entire expression are considered.
We estimate the model’s confidence level from the occurrences of confidence and uncertainty expressions throughout the reasoning process, following prior work (Liu et al., 2025) which measures LLM uncertainty by counting the frequency of epistemic markers. In Section 3, we use the category of expressions that appears more frequently near the end of the reasoning trace to determine whether the model is more confident or more uncertain. Building on this idea, Section 4 analyzes the trajectory of confidence and uncertainty signals over the entire reasoning sequence, captured by ConfDiff (Equation 2) and ConfCurv (Equation 3), to assess whether the model is more confident or more uncertain.
Appendix D System Prompt of BoostAbstention
We consider BoostAbstention (Kirichenko et al., 2025) which crafts a system prompt that encourages the model to abstain in designated scenarios, showing effectiveness for both reasoning and standard LLMs. For our experiments, we modify the system prompt to better match the abstention scenarios of questions beyond capability boundary:
Appendix E Additional Analysis of Capability Boundary Signals in Hidden States
In Section 5, we show that eventually solved and unsolved questions can be separated using signals extracted from the hidden states of the last input token. However, these signals may partially correlate with question difficulty rather than reflecting a purely intrinsic notion of capability boundary.
In this section, we do not aim to establish a clean separation between capability boundary signals and question difficulty. Instead, we provide additional analyses showing that the hidden state signals identified in our setup remain informative under stricter per-benchmark and cross-benchmark settings.
An intuition from human exams. One possible intuition comes from human test taking. In an exam with limited time and many difficult problems, skilled test-takers may quickly identify questions that are unlikely to be solved within the available budget and avoid wasting excessive time on them. This judgment may come either from recognizing that a question resembles problems they have repeatedly struggled with in the past (analogous to our Monitor strategy), or from taking a few initial steps and realizing that they lack confidence to proceed (analogous to our Monitor strategy). We present this analogy only as intuition, not as evidence of human-like metacognitive awareness in LRMs. Motivated by this intuition, we further conduct per-benchmark and cross-benchmark analyses as follows.
| Dataset | Model | LDA | LR |
| HLE | GPT-oss-20B | 89.3 | 90.3 |
| R1-Distill-Qwen3-8B | 94.9 | 96.4 | |
| R1-Distill-Qwen3-32B | 96.4 | 96.4 | |
| QwQ-32B | 93.4 | 96.4 | |
| GSM8k | GPT-oss-20B | 80.3 | 83.3 |
| R1-Distill-Qwen3-8B | 70.1 | 71.6 | |
| R1-Distill-Qwen3-32B | 85.2 | 89.0 | |
| QwQ-32B | 87.9 | 87.1 |
Per-benchmark test. We provide in Table 6 the accuracy of separating unsolvable questions from solvable ones using LDA and LR classifiers trained and tested on splits from the same dataset. Specifically, for HLE and GSM8K, we randomly split each dataset into an 80% training set and a 20% test set (with balanced solvable/unsolvable labels), train the linear classifier on the training set, and evaluate its classification accuracy on the test set.
The results show that linear classifiers can still separate solvable and unsolvable questions from hidden states, although with reduced accuracy compared to aggregated benchmarks setting (Table 2). This drop is small on HLE but larger on GSM8K. The reason is that HLE contains a larger proportion of unsolvable questions, enabling the linear classifiers to better learn the hidden state information associated with unsolvability. Consequently, these unsolvable questions in HLE can be effectively leveraged to develop models’ Hard Abstention (HA) capability.
Cross-benchmark test. In the above analysis, we think the unsolvable questions in HLE can be effectively leveraged to develop models’ Hard Abstention (HA) capability. Therefore, we train linear classifiers on HLE and evaluate models’ Hard Abstention capability on GSM8k and AIME to test whether this capability generalizes across benchmarks. As shown in Table 7, linear classifiers trained on HLE are able to identify nearly all unsolvable questions in GSM8K and AIME, achieving HA close to 100%.
| Model | GSM8k | AIME |
| GPT-oss-20B | 0 | 0 |
| +Monitor | 100 | 100 |
| R1-Distill-Qwen3-8B | 0 | 0 |
| +Monitor | 100 | 100 |
| R1-Distill-Qwen3-32B | 0 | 0 |
| +Monitor | 90.5 | 94.7 |
| QwQ-32B | 0 | 0 |
| +Monitor | 98.9 | 100 |
In contrast, training on GSM8K is insufficient for developing Hard Abstention capability, because GSM8K contains very few unsolvable questions. The linear classifiers cannot reliably learn the hidden state patterns associated with unsolvability.
Appendix F Ablation Study on Output Prefix
During our initial exploration, we experiment with several output prefixes and select the one that achieves the largest reduction in token usage. We conduct an ablation study comparing:
Initial prefix (our early attempt): Our goal is to prompt the model to acknowledge that it cannot solve an unsolvable question and then provide only a concise potential approach. However, we unexpectedly found that this cannot minimize token usage. The model’s instruction-following ability was insufficient for reliably producing only a concise approach, it often generated overly long responses. The initial prefix is shown below:
Emphasis prefix (final choice): Inspired by prior work (Zhang et al., 2024; Zhao et al., 2021) which show that the final sentence in the instruction strongly influences model behavior, we appended Step 1: at the end of the instruction to explicitly trigger step-by-step approach. Also, we emphasize in the end of the <think>...</think> part that I must give a concise outline to the user (less than 10 steps)!. This combination elicits short, structured responses and lead to the largest reduction in token usage. This is the output prefix we ultimately choose:
Token-budget prefix (following prior work): Following prior work (Han et al., 2025) which compresses unnecessary reasoning by including a reasonable token budget in the prompt, we replace less than 10 steps with an explicit budget constraint use less than 500 tokens. The token-budget prefix is:
| Output prefix | Token |
| No output prefix | 5096 |
| Initial | 4423 |
| Emphasis | 2227 |
| Token-budget | 2567 |
To avoid the confounding effect of incomplete answers, we set the context length to the maximum according to Huggingface model card.
Table 8 shows that all three designs reduce token usage, with the Emphasis prefix achieving the largest reduction. While it is possible that an even better prompt design exists, this does not affect our core conclusion: setting an output prefix enables the model to provide more efficient and more reliable responses on unsolvable questions.
Appendix G Analysis on Coding Task
| Model | Reasoning Expressions | Hidden States | ||
| ConfDiff | ConfCurv | LDA | LR | |
| GPT-oss-20B | 100 | 86.8 | 81.1 | 83.5 |
| R1-Distill-Qwen3-8B | 97.3 | 93.0 | 92.0 | 93.9 |
| R1-Distill-Qwen3-32B | 100 | 91.7 | 81.6 | 84.9 |
| QwQ-32B | 100 | 92.5 | 77.3 | 85.4 |
Since we capture unsolvability signals from reasoning expressions (Section 4) and hidden states (Section 5), without relying on task-specific heuristics, they might in principle generalize to other forms of reasoning tasks. We provide an analysis on a coding task in this section.
We conduct experiments on 1055 samples from LiveCodeBench (Jain et al., 2024). Table 9 shows the accuracy of separating solvable and unsolvable coding questions based on capability boundaries in reasoning expressions and hidden states.
We observe that coding tasks also exhibit clear capability boundaries in both reasoning expressions and hidden states. However, the accuracy from hidden states is lower than that from mathematical tasks. A plausible explanation is that the difficulty of coding problems is less explicitly signaled in the input itself. Unlike in math exams where humans can often immediately recognize “end-of-exam” questions that they are unlikely to solve, coding tasks usually require some degree of initial reasoning before the model can detect unsolvability. As a result, the unsolvability signal in the hidden states of the last input token is weaker, whereas the signal becomes more apparent in the early reasoning expressions.
Appendix H Capability Boundary in Hidden States
Figure 8 visualizes the classification of solvable and unsolvable questions of all LRMs, along with the capability boundary (blue line) determined by the decision boundary of linear classifier (e.g., LDA). We observe that the two classes of questions are clearly separated, under random train–test splits. These results indicate that the hidden states of the last input token contain useful signals related to the operational capability boundary, and is valid across different LRMs.
Appendix I Monitoring Strategies with Maximum Context Length
To evaluate whether our monitoring strategies (Monitor and Monitor) hold promise for optimizing long-context reasoning tasks (Ling et al., 2025; Kuratov et al., 2024), we provide the optimized token usage under extended token budget, set to the maximum according to HuggingFace model card (128k for GPT-oss-20B, 64k for R1-Distill-Qwen3-8B, and 32k for both R1-Distill-Qwen3-32B and QwQ-32B).
| Context Length | Token ↓ | Overflow (%) ↓ | |
| GPT-oss-20B | 128k | ||
| +Monitor | |||
| +Monitor | |||
| R1-Distill-Qwen3-8B | 64k | ||
| +Monitor | |||
| +Monitor | |||
| R1-Distill-Qwen-32B | 32k | ||
| +Monitor | |||
| +Monitor | |||
| QwQ-32B | 32k | ||
| +Monitor | |||
| +Monitor |
Under these settings, Table 10 shows that the models always produce complete answers (Overflow = 0), and our monitoring strategies still reduce token usage significantly as expected (Acc and HA are the same as under 2k or 4k context length).
Moreover, it is worth noting that our strategies are not constrained by context length:
Monitor continuously tracks the densities of confident and uncertain expressions throughout the reasoning process and can detect unsolvable questions at an early stage (Figure 3), well before the model reaches the end of reasoning.
Monitor identifies unsolvable questions directly from the hidden states of last input token, i.e., even before the model begins reasoning.
Because our strategies operate early in the reasoning trajectory, we intentionally used relatively short context length (e.g., 2k or 4k) in Section 6 to demonstrate their early detection capabilities.
Appendix J LLMs Usage
In this work, LLMs are used to polish part of the writing and to revise portions of the code. All core contributions are carried out by the authors, including the main ideas, literature review, experimental design and execution, figure and table preparation, and analysis.
Appendix K Unproductive Reasoning
Figure 9 and Figure 10 show representative unproductive reasoning phenomena in LRMs, including repetitive looping and error accumulation, which impair their capacity for sound and consistent reasoning in complex problem-solving scenarios. They respectively show how models get stuck in repetitive deductions (Figure 9) and how errors compound during formula application and result derivation (Figure 10). In mathematical problem reasoning, the former presents repeated deduction of the same uncertain conclusion, while the latter sees continuous error accumulation due to flawed initial methods.